OpenStudy (anonymous):

5m-2p=24 3m+2p=24 (Elimination using multiplication)

OpenStudy (anonymous):

it is already set up for elimination, add both equations you get 8m +0p = 48 8m = 48 you can solve from there?

OpenStudy (anonymous):

Yes I can! I just did not know to to set it up. Thank you!!

OpenStudy (anonymous):

there are other ways to solve it too

OpenStudy (anonymous):

that one was easy because it had a -2 and a+ 2 waiting for you to conveniently cancel out

OpenStudy (anonymous):

Ah, I see! It's a lot easier than I thought!

OpenStudy (anonymous):

do you want to know how to solve it a different way?

OpenStudy (anonymous):

Oh, and yes I do!

OpenStudy (anonymous):

5m-2p=24 3m+2p=24 both of these statements = 24 so each statement is = to the othe 5m-2p= 3m+2p gather the like variables 5m-2p= 3m+2p +2p +2p 5m =3m +4p -3m -3m 2m=4p divide both sides by 2 m=2p now in your original expressions everywhere you see an m, replace it with 2p this eliminates one variable.

OpenStudy (anonymous):

OMG, that is exactly what I needed! You are a fantastic person! I thank you for taking the time to help me! Have a good Friday!

OpenStudy (anonymous):

enjoy, and you are most welcome